# ContextGen: Contextual Layout Anchoring for Identity-Consistent Multi-Instance Generation

### v1

@@ 4.2 Baselines
We compare our method against a comprehensive set of state-of-the-art baselines across relevant domains. For layout-to-image generation, we include pioneering works such as LAMIC (Chen et al., 2025b) and MS-Diffusion (Wang et al., 2025). To evaluate spatial control, we also benchmark against CreatiLayout (Zhang et al., 2024), EliGen (Zhang et al., 2025a), MIGC (Zhou et al., 2024b), InstanceDiffusion (Wang et al., 2024), and GLIGEN (Li et al., 2023). In the domain of subject-driven generation, we benchmark against OmniGen2 (Wu et al., 2025b), DreamO (Mou et al., 2025), UNO (Wu et al., 2025c), XVerse (Chen et al., 2025a), and MIP-Adapter (Huang et al., 2024) to specifically assess identity preservation. For a cutting-edge benchmark, we highlight the latest proprietary models, including Nano Banana (formally named Gemini 2.5 Flash Image, Google’s latest multimodal model) (DeepMind, 2025) and GPT-4o-Image (OpenAI, ), as well as leading open-source models like Qwen-Image-Edit (Wu et al., 2025a) and FLUX.1-Kontext (Labs et al., 2025).
@@
Table 2: Quantitative results on COCO-MIG and LAMICBench++. Image-guided methods () use our pre-generated images by FLUX.1-Dev (Labs, 2024b)
Figure 6: Qualitative results on LayoutSam-Eval.
@@ 4.3 Comparison
Fig. 4 demonstrates our method’s superior performance in preserving both content and style across diverse scenarios. Our approach consistently maintains accurate object relationships and fine details where other methods fail - evident in the precise rendering of facial identities (old man’s wrinkles), object features (shape of the vase, appearance of piggy bank, color and texture of Sphynx cat).
@@
Qualitative analysis, as presented in Fig. 5, highlights two key capabilities of our method. First, our approach effectively handles instance overlap, a common challenge for existing methods which often leads to attribute leakage or instance missing/merging. Second, our method exhibits superior spatial layout control, allowing it to synthesize a coherent and well-structured image from source images that may lack consistency. Additionally, as demonstrated in Fig. 6, our method performs robustly on complex text prompts, accurately reflecting fine-grained textual details in the generated image while preserving precise layout control.
@@ 4.4 Ablation Study
We perform an ablation study to investigate the contribution of the ICA mechanism within our hierarchical attention architecture. We empirically divide the 57 DiT-blocks into three groups: FR-19 (first 19 blocks), MID-19 (middle 19 blocks), and BK-19 (last 19 blocks). The quantitative results on LAMIC-Bench++ are summarized in Tab. 3.
Prior work (Zhou et al., 2025) has demonstrated that MID-19 blocks have the most significant influence on instance-specific attributes. In alignment with this finding, our experiments confirm that applying the ICA mechanism selectively to the MID-19 blocks yields the highest average score of 64.66 and the best IDS score of 32.72. This configuration significantly outperforms the baseline that only uses the CLA mechanism, highlighting that targeted application of ICA is crucial for enhancing identity preservation and overall performance.
DPO Fine-tuning Analysis
To mitigate the model’s tendency to rigidly copy layout images while neglecting instance adaptation (e.g., posture, lighting), we employ Direct Preference Optimization (DPO) (Rafailov et al., 2024) with target images as preferred samples and layout images as less preferred. LoRA fine-tuning (Rank 256) is conducted with varying $\beta$ coefficients.
@@
1000 91.03 57.10 26.83 75.71 62.67
Tab. 4 and qualitative results in Sec. A.3 reveals three key findings:
  • Improved Composition: ITC and AES increase by $\uparrow$4.19–4.60% and $\uparrow$3.78–3.91% respectively across all $\beta$, demonstrating enhanced scene understanding.
  • Controlled Trade-off: IDS and IPS show moderate decreases ($\downarrow$6.49% and $\downarrow$1.07% at $\beta=1000$), with degradation scaling monotonically with $\beta$.
  • Optimal Configuration: $\beta=1000$ achieves the best balance (AVG 62.67), surpassing both non-DPO baseline (62.55) and other $\beta$ values.
This validates DPO’s ability to navigate the layout-adaptation trade-off.
@@ 5 Conclusion
In this work, we presented ContextGen, a novel framework for multi-instance generation that achieves precise control over layout and identity. Our approach is built on a unified token sequence that integrates text, layout, and multiple reference images, enabling a comprehensive understanding of the generation task. We introduced two core components: the Contextual Layout Anchoring (CLA) mechanism for enforcing robust spatial structure and the Identity Consistency Attention (ICA) mechanism for preserving fine-grained instance-specific attributes. Furthermore, our hierarchical attention architecture effectively leverages the intrinsic functional specialization of a Diffusion Transformer, with different layers dynamically attending to global and local contexts. To facilitate future research in this area, we created and will release the first large-scale, hierarchically-structured dataset, IMIG-100K, which we used to demonstrate our method’s superiority. Our extensive qualitative and quantitative evaluations show that ContextGen consistently outperforms state-of-the-art models, proving the efficacy of our design. We believe that this work provides a new foundation for the development of highly controllable and scalable multi-instance generation systems.
@@ A.1 Base Model Selection
In our framework, the selection of the base diffusion model is crucial for achieving high-fidelity multi-instance generation. We evaluated three variants of the FLUX family of models as potential backbones: FLUX.1-Dev (a general image generation model) (Labs, 2024b), FLUX.1-Fill (a local inpainting model) (Labs, 2024a), and FLUX.1-Kontext (an editing model) (Labs et al., 2025). While existing multi-subject-driven generation methods without layout control (Wu et al., 2025c; Hu et al., 2025) that do not rely on attention masks have successfully utilized FLUX.1-Dev, our experiments showed a significant limitation: without additional fine-tuning, FLUX.1-Dev failed to produce coherent images when an attention mask was applied. In contrast, both FLUX.1-Fill and FLUX.1-Kontext demonstrated the ability to generate images correctly with the attention mask. Among these two, FLUX.1-Kontext exhibited a noticeably superior capacity for identity preservation. Therefore, we chose FLUX.1-Kontext as the foundational model for our framework, leveraging its robust out-of-the-box performance with attention masking and its strong identity preservation capabilities.
@@ A.2 Details of Compositing Layout Image
A correct composition order is crucial for multi-instance synthesis, especially when handling occlusions and complex overlaps. We propose a dynamic sorting algorithm, Instance Layering Prioritization, which first handles explicit containment relationships by prioritizing instances whose masks are completely contained within another’s. For all other candidate instances, we use an innovative hybrid priority scoring system to simulate the natural layering of objects. We utilize a pre-processing step to obtain each instance’s precise effective area (Ravi et al., 2024). The priority score $P_{i}$ is calculated as:
@@
The proposed hybrid priority scoring system is designed to simulate general, high-probability composition orders for model training. The introduction of the random factor enhances data diversity and model robustness during training. Despite the supplementary Identity Consistency Anchoring (ICA) mechanism, the overlap relationships can still influence the final generated image. Thus, for inference, a user-provided layout offers a more direct and customized form of control.
A.3 Details of DPO Fine-tuning
(a) Results during the DPO fine-tuning process ($\beta=100$)
(b) Results at fine-tuning convergence with different DPO $\beta$
Figure 7: Results for the same input and fixed seed of DPO fine-tuning.
Appendix B More Experiment Details
@@ B.1 Experimental Details on LAMICBench++
As shown in Tab. 1, we benchmark our method against several strong baselines, including single-image-editing models (Wu et al., 2025a; Labs et al., 2025) and closed-source commercial models (OpenAI, 2024; DeepMind, 2025). Since these two categories of models require different input modalities and instruction methods, we prepared distinct inputs for a fair evaluation:

### v2

@@ 4.2 Baselines
We compare our method against a comprehensive set of state-of-the-art baselines across relevant domains. For layout-to-image generation, we include pioneering works such as LAMIC (chen2025lamic) and MS-Diffusion (wang2025msdiffusion). To evaluate spatial control, we also benchmark against CreatiLayout (zhang2024creatilayout), EliGen (zhang2025eligenentitylevelcontrolledimage), MIGC (zhou2024migc), 3DIS (zhou20243dis), InstanceDiffusion (wang2024instancediffusion), and GLIGEN (li2023gligen). In the domain of subject-driven generation, we benchmark against OmniGen2 (wu2025omnigen2), DreamO (mou2025dreamo), UNO (wu2025less), XVerse (chen2025xverse), and MIP-Adapter (huang2024resolvingmulticonditionconfusionfinetuningfree) to specifically assess identity preservation. For a cutting-edge benchmark, we highlight the latest proprietary models, including Nano Banana (formally named Gemini 2.5 Flash Image, Google’s latest multimodal model) (gemini25flash) and GPT-4o-Image (SimaAddendumTG), as well as leading open-source models like Qwen-Image-Edit (wu2025qwenimagetechnicalreport) and FLUX.1-Kontext (labs2025flux1kontextflowmatching).
@@ 4.3 Comparison
Fig. 4 demonstrates our method’s superior performance in preserving both content and style across diverse scenarios. Our approach consistently maintains accurate object relationships and fine details where other methods fail - evident in the precise rendering of facial identities (old man’s wrinkles), object features (shape of the vase, appearance of piggy bank, color and texture of Sphynx cat). More qualitative results are provided in Sec. D.1. Beyond fidelity, our model also exhibits high generative flexibility. A qualitative illustration of this adaptability is included in Sec. B.4, showcasing the model’s ability to modify subjects’ postures and attributes to comply with complex interactions given by text prompts (e.g., inter-subject interactions or dynamic scenes), thus proving it does not rigidly ”transfer” the references.
Table 2: Quantitative results on COCO-MIG and LayoutSAM-Eval. Image-guided methods () use our pre-generated images by FLUX.1-Dev (flux2024).
Figure 6: Qualitative results on LayoutSAM-Eval.
@@
Qualitative analysis, as presented in Fig. 5, highlights two key capabilities of our method. First, our approach effectively handles instance overlap, a common challenge for existing methods which often leads to attribute leakage or instance missing/merging. Second, our method exhibits superior spatial layout control, allowing it to synthesize a coherent and well-structured image from source images that may lack consistency. Additionally, as demonstrated in Fig. 6, our method performs robustly on complex text prompts, accurately reflecting fine-grained textual details in the generated image while preserving precise layout control. More qualitative results are provided in Secs. D.2 and D.3.
@@ 4.4 Ablation Study
We perform an ablation study to investigate the contribution of the ICA mechanism within our hierarchical attention architecture. We empirically divide the 57 DiT-blocks into three groups: FR-19 (first 19 blocks), MID-19 (middle 19 blocks), and BK-19 (last 19 blocks). The quantitative results on LAMICBench++ are summarized in Tab. 3.
Crucially, the baseline configuration (gray line in Tab. 3) that omits the CLA mechanism shows a substantial decline in performance across all metrics, reinforcing the indispensability of CLA for effective multi-instance generation.
Prior work (zhou2025dreamrenderer) has demonstrated that MID-19 blocks have the most significant influence on instance-specific attributes. In alignment with this finding, our experiments confirm that applying the ICA mechanism selectively to the MID-19 blocks yields the highest average score of 64.66 and the best IDS score of 32.72. Furthermore, the results in Tab. 3 suggest some potential functional focus across the FLUX-DiT layers, with a detailed analysis provided in Sec. B.6.
Importantly, while the value of the ICA may not be fully reflected by macro-level metrics alone, the ICA component is indispensable for preserving fine-grained identity details in scenes involving instance overlaps, as detailed in Sec. B.5.
@@
DPO Fine-tuning Analysis
To mitigate the model’s tendency to rigidly copy layout images while neglecting instance adaptation (e.g., posture, lighting), we employ Direct Preference Optimization (DPO) (rafailov2024directpreferenceoptimizationlanguage), utilizing target images as preferred samples and layout images as less preferred. We first conduct an ablation study on LoRA Rank 256 to determine the optimal $\beta$ coefficient (results summarized in Tab. 4). The optimal $\beta$ (1000) was subsequently applied to our final DPO fine-tuning, which uses the higher capacity LoRA Rank 512. The qualitative validation for the final model is provided in Sec. B.3.3.
Tab. 4 and further studies in Sec. B.3 reveals three key findings:
  • Improved Composition: ITC and AES metrics consistently show improvement across tested $\beta$ values, demonstrating DPO’s effectiveness in enhancing overall subject composition.
  • Controlled Trade-off: Identity metrics (IDS and IPS) exhibit a slight and controlled decrease, with degradation scaling monotonically with the regularization strength $\beta$.
  • Optimal Configuration: The setting of $\beta=1000$ yields the highest overall average (AVG 62.67). This configuration successfully navigates the fidelity-flexibility trade-off, outperforming both the non-DPO baseline (62.55) and other tested $\beta$ settings.
This validates DPO’s ability to navigate the fidelity-flexibility trade-off.
@@ 5 Conclusion
In this work, we presented ContextGen, a novel Diffusion Transformer-based framework for highly controllable multi-instance generation. The foundation of our approach is a unified contextual token sequence that integrates text, layout information, and multiple reference images, enabling a comprehensive understanding of the generation task. To tackle the twin challenges of precise spatial control and identity preservation, we introduced two dedicated mechanisms: the Contextual Layout Anchoring (CLA), which effectively enforces robust spatial structure using a composite layout image, and the Identity Consistency Attention (ICA), which ensures fine-grained instance-specific attribute preservation via a constrained attention strategy. Furthermore, our work involved an exploration of attention mechanism specialization across DiT layers, which informed the design of our hierarchical attention architecture. We also contributed the large-scale, hierarchically-structured dataset, IMIG-100K, complete with detailed layout and identity annotations, to accelerate future research in this domain. Extensive quantitative and qualitative evaluations confirm that ContextGen consistently outperforms state-of-the-art models, validating the efficacy and robustness of our proposed mechanisms. We believe this work provides a robust and scalable foundation for highly customizable image generation systems.
@@ A.1 Base Model Selection
In our framework, the selection of the base diffusion model is crucial for achieving high-fidelity multi-instance generation. We evaluated three variants of the FLUX family of models as potential backbones: FLUX.1-Dev (a general image generation model) (flux2024), FLUX.1-Fill (a local inpainting model) (flux1_fill_dev), and FLUX.1-Kontext (an editing model) (labs2025flux1kontextflowmatching). While existing multi-subject-driven generation methods without layout control (wu2025less; hu2025positionicunifiedpositionidentity) that do not rely on attention masks have successfully utilized FLUX.1-Dev, our experiments showed a significant limitation: without additional fine-tuning, FLUX.1-Dev failed to produce coherent images when an attention mask was applied. In contrast, both FLUX.1-Fill and FLUX.1-Kontext demonstrated the ability to generate images correctly with the attention mask. Among these two, FLUX.1-Kontext exhibited a noticeably superior capacity for identity preservation. Therefore, we chose FLUX.1-Kontext as the foundational model for our framework, leveraging its robust out-of-the-box performance with attention masking and its strong identity preservation capabilities.
@@ A.2 Details of Compositing Layout Image
A correct composition order is crucial for multi-instance synthesis, especially when handling occlusions and complex overlaps. We propose a dynamic sorting algorithm, Instance Layering Prioritization, which first handles explicit containment relationships by prioritizing instances whose masks are completely contained within another’s. For all other candidate instances, we use a hybrid priority scoring system to simulate the natural layering of objects. We utilize a pre-processing step to obtain each instance’s precise effective area (ravi2024sam2). The priority score $P_{i}$ is calculated as:
@@
The proposed hybrid priority scoring system is designed to simulate general, high-probability composition orders for model training. The introduction of the random factor enhances data diversity and model robustness during training. Despite the Identity Consistency Anchoring (ICA) mechanism, the overlap relationships can still influence the final generated image. Thus, for inference, a user-provided layout offers a more direct and customized form of control.
Appendix B More Details and Analysis on Experiments
@@ B.1 Experimental Details on LAMICBench++
As shown in Tab. 1, we benchmark our method against several strong baselines, including single-image-editing models (wu2025qwenimagetechnicalreport; labs2025flux1kontextflowmatching) and closed-source commercial models (openai2024gpt4ocard; gemini25flash). Since these two categories of models require different input modalities and instruction methods, we prepared distinct inputs for a fair evaluation: